OpenStudy (anonymous):

PLEASE HELP 6th GRADE EASY! Brahim completed 18 free throws and missed 12. Using experimental probability, how many free throws is he predicted to complete in the next 100 tries? A. 6 B. 36 C. 48 D. 60

OpenStudy (anonymous):

made 18 out of 30

OpenStudy (anonymous):

Hi satellite, after you finished here can you PLEASE help me?

OpenStudy (anonymous):

\[\frac{18}{30}=\frac{3}{5}\] or if you like \(.6\)

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

\[.6\times 100=60\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thank you! Thats what i thought i just was not sure!